EA010B

PI: Alef
Title: YZ CMi

Session: Oct 1996 (session 3)
Frequency: 3.6 cm
Mode: A
Obs time (UT): 305 EU 22:30(31/10)-09:30(01/11)
Stations: EbNtOnShSmUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 2

Comment


To give feedback on this experiment, click here

Experiment Feedback

Onsala (J. Conway, A. Polatidis at Mon Nov 4 11:25:12 EST 1996) Success
No known problems. Rain during the first part of the experiment.

Shanghai (Huang Xinyong at Tue Nov 5 03:56:27 EST 1996) Success
No problems. Clock offset 0.2 micro sec. Weather overcast.

JIVE ( at Tue Nov 5 05:37:20 EST 1996) Minor failures

HartRAO (Jonathan Quick at Tue Nov 5 07:40:12 EST 1996) Minor failures
Problems with setting up caused first pass of tape 1 to be scrambled. Otherwise no problems.

HartRAO (Jonathan Quick at Tue Nov 5 07:44:48 EST 1996) Minor failures
Problems with setting up caused first pass of tape 1 to be scrambled. Otherwise no problems.

Noto (Corrado Trigilio at Tue Nov 5 11:49:15 EST 1996) Success
No known problems. Clear sky.

Effelsberg (d graham at Wed Nov 6 09:32:09 EST 1996) Failure
no data, telescope would not drive

Urumqi (Liu Xiang at Tue Nov 19 03:47:47 EST 1996) Minor failures
UT23:55:50-00:24:46 and UT00:54:50-01:13:46, MK3 computer dead. Clock offset 0.4+/-0.5us.

MPIfR (Walter Alef at Tue Sep 23 17:37:37 METDST 1997) Minor failures
Cancelled due to loss of Effelsberg.